OpenStudy (anonymous):

What are the solutions for x in the proportion 4x-3 = 2x ---- ---- x 2 What are the solutions for x in the proportion 4x-3 = 2x ---- ---- x 2 @Mathematics

OpenStudy (eyust707):

cross multiply: 8x-6 = 2x^2 subtract 8x add 6 0=2x^2 -8x +6 can you do it from there?

OpenStudy (anonymous):

What are the solutions for x in the proportion 5x-6 = 2x ---- ---- x 2

OpenStudy (eyust707):

Factor out 2 0 = 2(x^2 -4x +3)

OpenStudy (eyust707):

now we can look at it. in order for that to be correct the right side needs to equal 0 well in order for the right side to be zero whats in side the ( ) needs to equal zero

OpenStudy (eyust707):

ive bet youve found the zeros of an equation like the one in the ( ) before =)

OpenStudy (anonymous):

Write a radical expression that has 4 terms (without variables), that are not in simplest form. Create this expression so that, once the terms are in simplest form, they can be combined to make an expression with one or two terms.

OpenStudy (eyust707):

do you know what a radical is?

OpenStudy (anonymous):

Nope.. Just need the answer to the question that.. it would be most helpful

OpenStudy (eyust707):

ill help you answer it if ya want. i try not to do peoples homework for them tho

OpenStudy (anonymous):

I understand.. but it would be quick and easier to just give the answer.. That's All

OpenStudy (eyust707):

yes yes it would.. but you wouldnt learn anything

OpenStudy (anonymous):

True.. Ok sbut get straight to the point. nothing extra

OpenStudy (eyust707):

ok so a radical is the name of a square rt sign.. \[\sqrt{}\] that thing

OpenStudy (eyust707):

so a radical expression is just numbers with that sign

OpenStudy (eyust707):

Do you know what a term is?

OpenStudy (anonymous):

I understand the radical sign and terms i just need to know how to solve this

OpenStudy (eyust707):

so basically what they want is ??? + ????+ ???? +???

OpenStudy (eyust707):

there needs to be a radical sign in it

OpenStudy (eyust707):

and each term needs to be more than just 1 number

OpenStudy (eyust707):

an example of this would be Sqrt(9)+ 2*3 + sqrt(25) - 3*1

OpenStudy (anonymous):

Could i use this example in this situation?

OpenStudy (eyust707):

no cause i did that one.. try and do your own

OpenStudy (anonymous):

Ok I got it.. but i have another Math Question

OpenStudy (anonymous):

What is the exact value of the expression \[5\sqrt{5}^+3 \sqrt{20}^-6\sqrt{3}\]

OpenStudy (eyust707):

ok so here we want to look at the \[3\sqrt{20}\]

OpenStudy (eyust707):

\[3\sqrt{20} \] also equals \[3\sqrt{5 * 4}\] which equals \[3\sqrt{5}\sqrt{4}\]

OpenStudy (eyust707):

and you know what the sqrt of 4 is